Find f(2) if f(x) = (x + 1)^2
Answer by mamiya(56)

find f(2) if f(x) = (x + 1)^2
you just have to plug 2 into f(x)
by plugging in 2 into f(x) , we get f(2) = (2+1)^2
= 3^2
= 9
so f(2) = 9